OpenStudy (anonymous):

find 5 rational numbers between -1 and 0

OpenStudy (zzr0ck3r):

\[\{\frac{-1}{6},\frac{-2}{6},\frac{-3}{6},\frac{-4}{6},\frac{-5}{6}\}\]

OpenStudy (anonymous):

thanks